OpenStudy (anonymous):

The doubling period of a bacterial population is 20 minutes. At time t = 100 minutes, the bacterial population was 80000. What was the initial population at time t = 0 ? Find the size of the bacterial population after 3 hours.

OpenStudy (dan815):

welcome to OS i will take good care of you

OpenStudy (anonymous):

can you help me with this problem?

OpenStudy (dan815):

yes

OpenStudy (dan815):

okay so every 20 mins the baceria doubles

OpenStudy (dan815):

0 to 20 to 40 to 60 to 80 to 100 at time 0, it was lets say k at 20 it was 2*k at 40 =2*2* K at 60 =2*2*2*k at 80 = 2*2*2*2*K at 100=2*2*2*2*2*K, or 2^5*K or 2^(100/20) * K we also know at 100 mins it was 80000 2^5*k=80,000 k=80,000/2^5

OpenStudy (dan815):

now for 300, there are a total of 300/20 = 30/2 = 15 periods so it will dobl 15 times 2^15*K=? we calculated k from earlier

OpenStudy (anonymous):

im not sure if I'm getting the 2nd part right….finding the size of bacterial population after 3 hours

OpenStudy (dan815):

wat did u get

OpenStudy (dan815):

oh oops

OpenStudy (dan815):

3 hours is not 15 doubling hehe

OpenStudy (anonymous):

haha

OpenStudy (dan815):

3 hours = 180 mins so 180/20=18/2=9 2^9*k=?

OpenStudy (anonymous):

are you able to help me with more? you're very helpful

OpenStudy (dan815):

dont worry there are lots of helpers

OpenStudy (dan815):

here

OpenStudy (anonymous):

A wooden artifact from an ancient tomb contains 45 percent of the carbon-14 that is present in living trees. How long ago, to the nearest year, was the artifact made? (The half-life of carbon-14 is 5730 years.)

OpenStudy (dan815):

make another question someone will help you soon if its not me

OpenStudy (dan815):

okay so

OpenStudy (anonymous):

Find the time required for an investment of 5000 dollars to grow to 6300 dollars at an interest rate of 7.5 percent per year, compounded quarterly.

OpenStudy (anonymous):

how about this one?